OpenStudy (anonymous):

Joe runs 8.75 times around a track in 1,125.803 seconds. If one lap around the track is 420.3 meters, which is the best estimate of the runner's average speed in meters per second (m/s)? 30 m/s 3 m/s 0.4 m/s 0.3 m/s

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (tkhunny):

Multiply by 1 (one)! \(\dfrac{8.75\;lap}{1125.803\;sec}\cdot\dfrac{420.3\;m}{1\;lap}\) I think it's done. Estimate away!

OpenStudy (anonymous):

haha i dont know

OpenStudy (anonymous):

i started with the left side then multiplied that times 420.3 and got 3.127125527

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

This is supposed to be an estimation problem. No fair calculating the exact value.

OpenStudy (anonymous):

im done with this